The treasurer for Pittsburgh Iron Works wishes to use financial futures to hedge her interest rate exposure. She will sell five Treasury futures contracts at $143,000 per contract. It is July and the contracts must be closed out in December of this year. Long-term interest rates are currently 11.30 percent. If they increase to 12.50 percent, assume the value of the contracts will go down by 5 percent. Also if interest rates do increase by 1.2 percent, assume the firm will have additional interest expense on its business loans and other commitments of $46,000. This expense, of course, will be separate from the futures contracts. a. What will be the profit or loss on the futures contract if interest rates go to 12.50 percent by December when the contract is closed out? (Input the amount as a positive value.) on futures contracts $ b-1. After considering the hedging, what is the net cost to the firm of the increased interest expense of $46,000? Net cost $ b-2. What percent of this $46,000 cost did the treasurer effectively hedge away? (Input your answer as a percent rounded to 2 decimal places.) Percentage hedged away % c. Indicate whether there would be a profit or loss on the futures contracts if interest rates went down. Loss Profit HintsReferenceseBook & Resources WorksheetDifficulty: ChallengeLearning Objective: 08-05 Hedging may be used to offset the risk of interest rates rising. Check my work
Answer #1

A)

Profit at 12.5% int rate/ Future Contract = 143000 - (143000 - 5% of 143000)

                                                              = 7150

Total profit for 5 contracts = 7150 * 5

                                         = 35750.00

B-1)

Net cost after hedging = 46000 - 35750

                                   = 10250

B-2)

% hedged = 35750/46000 * 100

                  = 77.72